OpenStudy (anonymous):

In a geometric sequence, a5=8 and a7=32. use the geometric mean to find the value of a6.

OpenStudy (anonymous):

since a5, a6 and a7 are geometric progression therefore \[a_6^2= a_5 \times a_7 \] \[a_6^2= 8 \times 32 =256 \rightarrow a_6= \sqrt{256}\] \[a_6= \pm16\] since 8 and 32 are positive numbers therefore \[a_6= 16\]

OpenStudy (anonymous):

thank you! @dpasingh